### v1

@@ 2.2 Complexity of domain and range
The following hold for $f:\subseteq 2^{\omega}\to 2^{\omega}$:
@@
    if $x\in\Delta^{1}_{1}$ there is a partial computable $f$ with domain $\{z\}$ and $x\leq_{T}z$
@@
    the domain of every partial computable $f$ contains a real in $\Delta^{1}_{1}$.
@@
For each $x\in\Delta^{1}_{1}$ there is an ordinal $\alpha<\omega_{1}^{\rm CK}$ such that $x\leq_{T}\emptyset^{(\alpha)}$. By [7, Theorem 2.1.4] $\{\emptyset^{(\alpha)}\}$ is $\Pi^{0}_{2}$. So (i) holds by Proposition 2.1 (iii). By the Gandy Basis Theorem [24, III.1.5] every nonempty $\Sigma^{1}_{1}$ class has a $\Delta^{1}_{1}$ member. So (ii) holds by Proposition 2.1 (i). ∎
@@ 3.1 Inversion without random oracles
There is a partial computable injection $f:\subseteq 2^{\omega}\to 2^{\omega}$ such that no arithmetically definable $g$ can invert $f$ on any real.
@@
Any $g$ that inverts $f$ on $y=0^{\omega}$ defines $\emptyset^{(\omega)}$ which is not arithmetical. ∎

### v2

@@ 2.2 Complexity of domain and range
The following hold for partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$:
@@
    if $x\in\Delta^{1}_{1}$ there is $f$ with domain $\mathsf{dom}(f)=\{z\}$ and $x<_{T}z$
@@
    there is $f$ with uncountable $\mathsf{dom}(f)$ and $\mathsf{dom}(f)\cap\Delta^{1}_{1}=\emptyset$.
@@
For each $x\in\Delta^{1}_{1}$ there is an ordinal $\alpha<\omega_{1}^{\rm CK}$ such that $x<_{T}\emptyset^{(\alpha)}$. By [7, Theorem 2.1.4] $\{\emptyset^{(\alpha)}\}$ is $\Pi^{0}_{2}$. So (i) holds by Proposition 2.1 (iii). Let
be a computable injection from $\omega^{\omega}$ into $2^{\omega}$. By [7, Proposition 2.5.2] there is a computable tree $T\subseteq\omega^{\omega}$ such that $[T]$ is uncountable and $[T]\cap\Delta^{1}_{1}=\emptyset$.
Every $x\in\pi([T])$ computes a member of $[T]$ so $\pi([T])\cap\Delta^{1}_{1}=\emptyset$. Since the image $\pi([T])$ is a $\Pi^{0}_{2}$ subset of $2^{\omega}$ we get (ii) by Proposition 2.1 (iii). ∎
@@ 3.1 Inversion without random oracles
There are partial computable $f,h:\subseteq 2^{\omega}\to 2^{\omega}$ such that
  1. (i)
    $f$ is injective and no arithmetically definable $g$ can invert $f$ on any real
  2. (ii)
    $\mathsf{dom}(h)$ is uncountable and no $g\in\Delta^{1}_{1}$ can invert $h$ on any real.
@@
Any $g$ that inverts $f$ on $y=0^{\omega}$ defines $\emptyset^{(\omega)}$ which is not arithmetical.
By Proposition 2.3 (ii) let $P$ be an uncountable $\Pi^{0}_{2}$ class with no $\Delta^{1}_{1}$ members. The restriction of $x\mapsto 0^{\omega}$ to $P$ meets the requirements of (ii). ∎